# Spherical Antenna Arrays: Principles, Applications, and Research Directions

### 03 Apr 2026 v1

@@ I-A Bottlenecks of Linear and Planar Arrays
Traditional ULAs and UPAs exhibit inherent limitations that severely constrain system performance in practical wireless communication applications. ULAs support beam scanning in only one dimension [2], resulting in limited angular resolution and narrow spatial coverage, which renders them incapable of addressing multi-directional or dynamically varying communication targets in 3D space. Although UPAs enable two-dimensional beam steering, they suffer from substantial performance degration, with angular resolution decreasing significantly as the target departs from the boresight direction [3, 4]. In 3D scenarios such as low-altitude unmanned aerial vehicle (UAV) swarms, this often leads to issues including signal outage, mutual coupling, and coverage holes. Furthermore, improving the resolution of conventional arrays typically requires to increase the array aperture and the number of elements, which in turn substantially increases the complexity of RF chains and phase shifters, thereby increasing system power consumption, cost, and physical size [5]. Such trade-offs conflict with the application requirements of miniaturized platforms such as low-altitude terminals and satellites. Collectively, these limitations render traditional arrays unsuitable to meet the demands of future wireless communication applications, including ISAC, UAV swarm networking, and satellite communications [6, 7].
@@ I-B Advantages of Spherical Antenna Arrays
An SAA is a type of 3D antenna array in which elements are uniformly distributed on a spherical surface, forming a symmetric spatial structure. Unlike conventional ULAs or UPAs confined to one or two spatial dimensions, SAAs leverage the inherent geometric properties of the sphere to enable multi-directional signal interaction, as shown in Fig. 1. Owing to their 3D spherical geometry, SAAs offer a range of distinct performance advantages over traditional linear and planar arrays, making them as a promising solution for overcoming the limitations of conventional array architectures.
@@ I-C Typical Structures of Spherical Antenna Arrays
Several representative geometric structures of SAAs are illustrated in Fig. 2.
The equal angular ring structure arranges elements along multiple parallel circular rings at equal angular intervals. As the most classical configuration, it has the benefits of geometric regularity and analytical tractability [11].
The polyhedral projection structure maps the vertices of a subdivided regular polyhedron onto the spherical surface, yielding highly uniform element spacing and excellent spherical symmetry. This structure is particularly suitable for high-performance massive MIMO and spatial signal processing systems [8].
The spherical spiral structure places elements along a continuous spiral curve on the spherical surface, avoiding periodic structural repetition and enabling quasi-continuous spatial coverage. This configuration is advantageous for broadband beamforming and high-resolution direction estimation [12].
@@ II-B Satellite Communications
Unlike planar arrays that suffer gain degradation at large scan angles, the spherical geometry maintains consistent gain in all directions. A single SAA-equipped BS can track satellites across Earth orbits without physical rotation, avoiding the link outage. For inter-satellite links, the ability to simultaneously form multiple beams enables a single array to maintain connections with multiple satellites across the entire satellite network.
@@ III-A Angular Beam Pattern
The angular beam patterns of the SAA and UPA are illustrated in Fig. 5, with key simulation parameters detailed as follows. The radius of the SAA is set to $R=0.5\,\text{m}$, while the element spacing of the UPA is $d=\lambda/2$. Both arrays are configured with $N=100$ antenna elements, operating at a wavelength of $\lambda=0.01\,\text{m}$. The focal distance is fixed at $r_{0}=30\,\text{m}$. Three focal points are adopted for performance evaluation, where the azimuth angle is fixed at $\phi_{0}=\pi/3$ ($60^{\circ}$) and the elevation angles are set to $\theta_{0}=0$, $\pi/3$, and $2\pi/3$ (corresponding to $0^{\circ}$, $60^{\circ}$, and $120^{\circ}$), respectively. The beam patterns are assessed over the elevation range $\theta\in[0,\pi]$ (i.e., $0^{\circ}$ to $180^{\circ}$), with a total of 181 sampling points used for discrete evaluation.
Figure 5: Angular Beam Patterns with Three Focal Points
@@ III-B Distance Beam Pattern
Fig. 6 shows the normalized beam power distribution of SAAs with different radii along the distance dimension. It can be seen that, similar to the UPA, the SAA also has near-field distance-focusing capability. With the increase of the SAA radius $R$, the distance-focusing effect is significantly enhanced: the beam peak becomes sharper, the side-lobe is lower, and the energy concentration is higher. In particular, when the radius is $R=2\,\text{m}$, the SAA can achieve precise focusing at the target distance of 30 m, with strong range selection ability and high focusing resolution.
@@ V Conclusion
This article has demonstrated that SAAs offer unique advantages over conventional ULAs and UPAs, including 360° full-space coverage, isotropic angular resolution, and superior near-field focusing capability. A comparative case study confirms that SAAs maintain stable beamforming performance across all directions, while planar arrays fail to cover the backward hemisphere. These features make SAAs attractive for emerging 6G applications such as integrated sensing and communication, satellite communications, and low-altitude economy. Despite these benefits, practical deployment faces critical challenges in channel modeling, estimation, codebook design, beam training, and localization. Addressing these issues through interdisciplinary research will position SAAs as a key enabling technology for future wireless communication systems.

### 08 Apr 2026 v2

# Spherical Antenna Arrays for Future Communications: Principles, Applications, and Research Directions

@@ I-A Bottlenecks of Linear and Planar Arrays
Traditional ULAs and UPAs exhibit inherent limitations that severely constrain system performance in practical wireless communication applications. ULAs support beam scanning in only one dimension [2], resulting in limited angular resolution and narrow spatial coverage, which renders them incapable of addressing multi-directional or dynamically varying communication targets in 3D space. Although UPAs enable two-dimensional beam steering, they suffer from substantial performance degration, with angular resolution decreasing significantly as the target departs from the boresight direction [3, 4, 5]. In 3D scenarios such as low-altitude unmanned aerial vehicle (UAV) swarms, this often leads to issues including signal outage, mutual coupling, and coverage holes. Furthermore, improving the resolution of conventional arrays typically requires to increase the array aperture and the number of elements, which in turn substantially increases the complexity of radio frequency (RF) chains and phase shifters, thereby increasing system power consumption, cost, and physical size [6]. Such trade-offs conflict with the application requirements of miniaturized platforms such as low-altitude terminals and satellites. Collectively, these limitations render traditional arrays unsuitable to meet the demands of future wireless communication applications, including ISAC, UAV swarm networking, and satellite communications [7, 8].
@@ I-B Advantages of Spherical Antenna Arrays
An SAA is a type of 3D antenna array in which elements are uniformly distributed on a spherical surface, forming a symmetric spatial structure. Unlike conventional ULAs or UPAs confined to one or two spatial dimensions, SAAs leverage the inherent geometric properties of the sphere to enable multi-directional signal interaction, as shown in Fig. 1. Owing to their 3D spherical geometry, SAAs offer a range of distinct performance advantages over traditional linear and planar arrays, making them a promising solution for overcoming the limitations of conventional array architectures.
@@ I-C Typical Structures of Spherical Antenna Arrays
Several representative geometric structures of SAAs are illustrated in Fig. 2. The equal angular ring structure arranges elements along multiple parallel circular rings at equal angular intervals. As the most classical configuration, it has the benefits of geometric regularity and analytical tractability [12]. The polyhedral projection structure maps the vertices of a subdivided regular polyhedron onto the spherical surface, yielding highly uniform element spacing and excellent spherical symmetry. This structure is particularly suitable for high-performance massive MIMO and spatial signal processing systems [9]. The spherical spiral structure places elements along a continuous spiral curve on the spherical surface, avoiding periodic structural repetition and enabling quasi-continuous spatial coverage. This configuration is advantageous for broadband beamforming and high-resolution direction estimation [13].
@@ II-B Satellite Communications
Unlike planar arrays that suffer from gain degradation at large scan angles, the spherical geometry maintains consistent gain in all directions. A single SAA-equipped BS can track satellites across Earth orbits without physical rotation, avoiding the link outage. For inter-satellite links, the ability to simultaneously form multiple beams enables a single array to maintain connections with multiple satellites across the entire satellite network.
@@ III-A Angular Beam Pattern
The angular beam patterns of the SAA and UPA are illustrated in Fig. 5, with key simulation parameters detailed as follows. The radius of the SAA is set to $R=0.5\,\text{m}$, while the element spacing of the UPA is $d=\lambda/2$. Both arrays are configured with $N=100$ antenna elements, operating at a wavelength of $\lambda=0.01\,\text{m}$. The focal distance is fixed at $r_{0}=30\,\text{m}$. Eight focal points are adopted for performance evaluation, with azimuth and elevation angles specified as: $(\phi,\theta)$ == $(\pi/6,\pi/6)$, $(5\pi/6,\pi/3)$, $(\pi/3,\pi/4)$, $(3\pi/4,2\pi/3)$, $(\pi/4,3\pi/4)$, $(2\pi/3,5\pi/6)$, $(\pi/3,0)$, $(\pi/3,\pi)$. The beam patterns are assessed over the full angular range, i.e., $\theta\in[0,\pi]$ (elevation) and $\phi\in[0,2\pi]$ (azimuth), with $181\times 181$ sampling points used for discrete evaluation.
Figure 5: Angular Beam Patterns with Eight Focal Points
@@ III-B Distance Beam Pattern
Fig. 6 shows the normalized beam power distribution of SAAs with different radii along the distance dimension. It is observed that, similarly to the UPA, the SAA also possesses near-field distance-focusing capability. As the radius $R$ of the SAA increases, the distance-focusing effect is significantly enhanced: the main beam becomes sharper, the side-lobe levels decrease, and the energy concentration improves. Notably, when the radius reaches $R=2\,\text{m}$, the SAA achieves precise focusing at the target distance of 30 m, exhibiting strong range selectivity and high focusing resolution. It should also be noted that the distance-focusing capability of SAAs can be further improved by employing higher operating frequencies, such as those in the terahertz (THz) bands. In fact, the distance-focusing capability of an SAA is directly related to the ratio of the array radius to the operating wavelength: a larger ratio yields a stronger distance-focusing capability.
@@ V Conclusion
This article has demonstrated that SAAs offer unique advantages over conventional ULAs and UPAs, including 360° full-space coverage, isotropic angular resolution, and superior near-field focusing capability. A comparative case study confirmed that SAAs maintain stable beamforming performance across all directions, while planar arrays fail to cover the backward hemisphere. These features make SAAs attractive for emerging 6G applications such as integrated sensing and communication, satellite communications, and low-altitude economy. Despite these benefits, practical deployment faces critical challenges in channel modeling, estimation, codebook design, beam training, and localization. Addressing these issues through interdisciplinary research will position SAAs as a key enabling technology for future wireless communication systems.